OpenStudy (anonymous):

olve one-seventh plus five-sixths. Then use estimation to check for reasonableness. Show your w

OpenStudy (anonymous):

@Asevilla5 help plz

OpenStudy (anonymous):

To add or subtract fractions you need a common denominator. The LCD of 7 and 6 is 42. 1/7=6/42. 5/6=35/42. 6/42+35/42 = 41/42. Therefore, the answer is 41/42.

OpenStudy (anonymous):

Using estimation to check the answer, you can round 1/7 to zero and 5/6 to one. Zero plus one is one, and you can round 41/42 to one.